# Improvement of Spiking Neural Network with Bit Planes and Color Models

### v3

@@ Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the SNN for image classification through a new coding method that exploits bit plane representation. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets of grayscale images and color images. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more effective SNNs models in future researches and applications. Source code for our experiments is available at https://github.com/luutn2002/bit-plane-snn
@@ I Introduction
$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of $colormodel$
$colormodel$ A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
$mod$ Modulo division
@@ II-A Overview of SNNs
SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm. Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18]. This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19]. While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20]. Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21]. Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9] or ANN-to-SNN conversion approaches [22]. Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
@@ III-A Bit planes in deep learning
Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored. As far as we concern, our research is first to explore application of bit planes on SNNs spike coding process.
@@ IV Proposed Method
function BitplaneEncode($X,X_{max}$, $colormodel$)
     $X=convertColor(X,colormodel)$triangleright Convert $X$ to specified color model
     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
     $L\leftarrow[\text{ }]$triangleright Empty list initialization
@@
              $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane to $L$
              $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of input image tensor and floor the result
@@
     return $concat(L)$triangleright Concatenate all stored sub tensors in $L$ to a single tensor
@@ V Theoretical basis
where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction.
@@
$\delta(\mathbf{x})=h(\mathbf{x})g_{\mathbf{w}}(\mathbf{x})$
@@
To improve the SNR, we can either increase signal term $Sig(\delta(\mathbf{x}))$ or reduce noise term $Noi(\delta(\mathbf{x}))$. As shown in [11], the norm $\left\|\delta({\mathbf{x}})\right\|$ is positively correlated with timestep $T$:
$\left\|\delta({\mathbf{x}})\right\|=\left\|h^{\prime}({\mathbf{x}})\right\|\cdot\left\|\frac{\partial\ell}{\partial O^{l+k-1}}\right\|\cdot\left\|\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\right\|$ (11)
$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)=\sqrt{NT\alpha}$
$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]}{N\,T}$
where the equation denote the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ with target function gradient $h^{\prime}({\mathbf{x}})$, previous blocks’ gradient $\frac{\partial L}{\partial O^{l+k-1}}$ of a $k$ sequential blocks $l$ layers Spiking-ResNet architecture[11] and Heaviside function’s gradient $\Theta^{\prime}$ while transmitting signal $S$ of previous layer output $O$ at firing rate $\Phi$ using $N$ neurons with threshold voltage $V_{th}$.
Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ and firing rate $\Phi$ would increase as:
[…]
$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\Theta^{\prime}\left(B^{l+i}-V_{th}\right)=\sqrt{N(T+N_{bit})\alpha}$
$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]+\sum_{j=0}^{N-1}\sum_{t=0}^{N_{bit}-1}B^{l}_{j}[t]}{N\,(T+N_{bit})}$
and improving the chance of obtaining higher $SNR(\delta(\mathbf{x}))$.
@@ VI-A Experimental settings
$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\geq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (13)
To enable efficient training, surrogate gradient methods were applied to all IF neurons. We utilized the arctangent surrogate gradient function $S(U)$[11, 10]. Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[45, 50, 51], computed using mean reduction across mini-batches.
@@ VI-B2 Comparison of SNN variants
TABLE V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants. Performance gain when compared with the baseline is shown in green.
@@ VI-C Discussion
TABLE VIII: Comparison of model training times between baseline and proposed method using different color models. The values are average training times per epoch.
Model Variants Processing time by dataset (ms)
CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average
Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26
Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%)
Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%)
Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%)
Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%)
Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%)
Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%)
Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
@@
Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs. Despite the significant performance gains, the added computational complexity is a limitation of our method. As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications. Also, another limitation is that, similar to recent studies [26, 11], the experiments in our study are still based on software simulations. Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions. First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs, especially for different machine learning tasks. Second, it is obvious that the choice of color model may depend on the task in question, such as medical image analysis or traffic object recognition. So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks. In addition, the proposed method should be implemented on some hardware platforms and tested in practical scenarios.

### v4

@@ Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the SNN for image classification through a new coding method that exploits bit plane representation. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets of grayscale images and color images. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more effective SNNs models in future researches and applications.
@@ I Introduction
$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of colormodelcolormodel
colormodelcolormodel A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
modmod Modulo division
@@ II-A Overview of SNNs
SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm. Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18]. This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19]. While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20]. Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21]. Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9, 11, 22], hybrid SNN training [23] or ANN-to-SNN conversion approaches [24]. Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
@@ III-A Bit planes in deep learning
Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored. As far as we are concerned, our research is first to explore application of bit planes on SNNs spike coding process.
@@ IV Proposed Method
function BitplaneEncode($X,X_{max}$, colormodelcolormodel)
  $X=convertColor(X,colormodel)$ $\triangleright$ Convert $X$ to specified color model
  $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$ $\triangleright$ Get number of bit required to encode image from highest possible value $X$
  $L\leftarrow[\text{ }]$ $\triangleright$ Empty list initialization
@@
     $L.insert(X\mod{2})$ $\triangleright$ Element-wise modulo of tensor, then store the bit plane to $L$
     $X\leftarrow\left\lfloor X/2\right\rfloor$ $\triangleright$ Element-wise division of input image tensor and floor the result
@@
  return $concat(L)$ $\triangleright$ Concatenate all stored sub tensors in $L$ to a single tensor
@@ V Theoretical basis
where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction (basis for this statement can be found ar Appendix A).
@@
$\delta(\mathbf{x})=h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x})$
@@
Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, given certain condition satisfied, we can have:
$\mathrm{sup}\left(SNR\left(\delta^{bit}(\mathbf{x})\right)\right)\geq\mathrm{sup}\left(SNR\left(\delta(\mathbf{x})\right)\right),$ (11)
$\mathrm{sup}\left(SNR(\delta^{bit}(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}},$
$\mathrm{sup}\left(SNR(\delta(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}$
and improving the chance of obtaining higher SNR (more background details and proof for Equation 11 can be found at Appendix B).
@@ VI-A Experimental settings
$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
To enable efficient training, surrogate gradient methods were applied to all IF neurons. We utilized the arctangent surrogate gradient function S(U)S(U)[11, 10]. Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[47, 52, 53], computed using mean reduction across mini-batches.
@@ VI-B2 Comparison of SNN variants
Table V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants. Performance gain when compared with the baseline is shown in green and best performing model on average is denoted in bold.
@@ VI-C Discussion
Table VIII: Comparison of model total forward and backward propagation time between baseline and proposed method using different color models. The values are average training times per epoch per images batch of 16
. Model Variants Processing time by dataset (ms) CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26 Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%) Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%) Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%) Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%) Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%) Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%) Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
@@
Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs. Despite the significant performance gains, the added computational complexity is a limitation of our method. As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications. Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[3]. Empirical evaluation of processing speed and utilized memory could varied across datasets, attributed to factors such as input resolution, hardware specifications, and software library versions. Also, another limitation is that, similar to recent studies [28, 11], the experiments in our study are still based on software simulations. Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions. First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[63]. Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[64] , object recognition[65] and knowledge distillation [66]. So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks. Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[6] and tested in practical scenarios such as low power edge devices[67].